OpenStudy (anonymous):

When completely factored, 3x2 - 48 equals

OpenStudy (anonymous):

3x^2 - 48 = 0 divide both sides by 3 x^2 - 16 = 0 (x + a)(x + b) Variables 'a' and 'b' should be equal in value, but opposite in polarity. Also, the product of 'a' and 'b' must be -16 The possible factors of -16 include -4 and 4 Therefore, if a = -4, then b = 4, or if a = 4, then b = -4. Assuming that a = -4 and b = 4, we have (x - 4)(x + 4) because (x - 4)(x + 4) = x^2 - 4x + 4x - 16 = x^2 - 16 Assuming that a = 4 and b = -4, we have (x + 4)(x - 4) because (x + 4)(x - 4) = x^2 + 4x - 4x - 16 = x^2 - 16 So the factors are (x + 4) and (x - 4). Therefore, x = -4 or x = 4

OpenStudy (anonymous):

8

OpenStudy (anonymous):

how is it 8 ? o_O